What is the least five-digit number which is divisible by 8, 9, and 12?
Answer by stanbon(75887)

What is the least five-digit number which is divisible by 8, 9, and 12?
-----
====================
8 = 2^3
9 = 3^2
12 = 2^2*3
----
lcm = 2^3*3^2 = 72
-----
least to have 5 digits::: 72*2^8 = 18432
=================
Cheers,
Stan H.